OpenStudy (anonymous):

ln x + ln(x+3) = ln10

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

for left side, it can simplified b using the property of log ln x + ln(x+3) = ln x(x+3) now u have ln x(x+3) = ln 10 cancel out the ln from both sides, giving us x(x+3) = 10 x^2 + 3x - 10 = 0 solve for x (hint : this can be factored)

OpenStudy (anonymous):

thank you can you help me with another one again?

OpenStudy (anonymous):

:)

OpenStudy (raden):

i have to go to school now, if not i can come late :)

OpenStudy (anonymous):

np ^^

OpenStudy (raden):

see u ;)

OpenStudy (anonymous):

m'kay :)

OpenStudy (raden):

well, i want continue for the rest x^2 + 3x - 10 = 0 (x+5)(x-2) = 0 the zeroes satisfies if x+5 = 0 or x-2 = 0 we get x = -5 or x = 2

OpenStudy (raden):

but we have to check of them again. we subtitute them to the original equation : ln(x) + ln(x+3) = 10 if x=-5 we have ln(-5) + ln(-5+3) = 10 actually, it does not can be a solution for x = -5, because the log(x) would be undefined. if x = 2 we have ln(2) + ln(2+3) = ln 10 it can be a solution because ln(x) defined for x = 2 thus, the solution is only x = 2